Consider the following hypothetical chemical react
Answer #1

If same Heat of reaction

assume

enthalpy = Hrxn

same kinetics so ignore that, that has nothing to do with equilibrium

entropy = more mol of gas, more entropy generation

Then, according to Gibbs

G = H - TS

if in all cases H is the same, we might ignore, let us concentrate in TS term

assume same temperature so

only S is a factor

in order to G to be in forward reaction, it mus be negative i.e. G < 0

that is

G = H - TS = S (according to our study)

so

S > 0

S generation must be HIGH

so

A)

3 mol of gas turn to 2 mol o fgas, decrease

B)

3 mol of gas = 4 mol of gas, increase

3 mol of gas = 3 mol of gas, same

Therefore, choose B, since mol of gas increase, that is, entropy increase, therefore -TS increases and G is likely to be forward